Suppose you wish to determine the proportion of college students nationally who receive some form of financial aid. You want to be 98% confident of your results and have a maximum error of 5%. Calculate the minimum sample size that you must have to meet these requirements, given that the financial aid office at a local institution estimates the percentage to be 78%.
Answer #1

Solution :

Given that,

= 0.78

1 - = 0.22

margin of error = E = 0.05

Z/2 = 2.326

sample size = n = (Z / 2 / E)2 * * (1 - )

= (2.326 / 0.05)2 * 0.78 * 0.22

= 372

sample size = n = 372